# Efficient Artifacts Removal for Adaptive Deep Brain Stimulation and a Temporal Event Localization Analysis

### v1

@@ Abstract.
Objective. Adaptive deep brain stimulation (aDBS) leverages symptom-related biomarkers to deliver personalized neuromodulation therapy, with the potential to improve treatment efficacy and reduce power consumption compared to conventional DBS. However, stimulation-induced signal contamination remains a major technical barrier to advance its clinical application. Approach. Existing artifact removal strategies, both front-end and back-end, face trade-offs between artifact suppression and algorithmic flexibility. Among back-end algorithms, Shrinkage and Manifold-based Artifact Removal using Template Adaptation (SMARTA) has shown promising performance in mitigating stimulus artifacts with minimal distortion to local field potentials (LFPs), but its high computational demand and incapability of handling transient direct current (DC) artifacts limits its use in real-time applications. To address this, we developed SMARTA+, a computationally efficient extension of SMARTA capable of suppressing both stimulus artifacts and DC transient artifacts while supporting flexible algorithmic design. Main results. We evaluated SMARTA+ using semi-real aDBS data and real data from Parkinson’s disease patients. Compared to SMARTA and other established artifact suppression methods, SMARTA+ achieved comparable or superior artifact removal while significantly reducing computation time. It preserved the spectral and temporal structure, ranging from beta band to high frequency oscillation, of the underlying LFPs and demonstrated robustness across a variety of simulated stimulation protocols. Furthermore, the temporal event localization analysis show that SMARTA+ can more accurately determine the beta burst events compared with existing tools. Significance. By enhancing artifact suppression and improving computational efficiency, we show that SMARTA+ is potential for advancing real-time, closed-loop aDBS systems for neuromodulation therapies across diverse neurological disorders.
@@ 1. Introduction
Deep brain stimulation (DBS) is a widely used and effective therapy for advanced Parkinson’s disease (PD) [21, 3, 20, 8]. By delivering electrical stimulation through implanted electrodes in the subthalamic nucleus (STN), DBS can alleviate motor impairments and improve patients’ quality of life. Additionally, local field potentials (LFPs) can be recorded from the same electrodes during stimulation, providing real-time insights into PD-related neural activity. Biomarkers derived from LFP enable adaptive deep brain stimulation (aDBS), where stimulation parameters are dynamically adjusted to respond to fluctuations in disease state [23, 2]. The clinical benefits of aDBS have been demonstrated in PD, with improvements observed in motor symptoms such as tremor [40] and bradykinesia [22]. Beyond its application in PD, aDBS has also shown promise in treating essential tremor [41, 29] and epilepsy [19, 33]. These studies suggest that aDBS may be broadly applicable to various movement disorders, offering a more efficient and patient-specific therapeutic strategy.
@@
Various techniques, spanning both front-end and back-end processing, have been developed to mitigate stimulation artifacts in aDBS recordings. A basic front-end approach is common-mode rejection, where two sensing contacts adjacent to the stimulation electrode are used to obtain bipolar LFPs by signal subtraction [23, 2]. Ideally, when contacts are symmetrically positioned around the stimulation site, artifacts cancel out. However, impedance mismatches and tissue asymmetries often leave residual artifacts and DC transient distortions [1]. To further reduce artifacts, blanking techniques can be applied at the front-end [32] or back-end [14], temporarily disabling signal acquisition during stimulation pulses. While effective at suppressing high-amplitude artifacts, blanking also removes underlying neural signals and fails to address transient DC artifacts that persist beyond the stimulation period. When the frequency band of interest is known in advance, such as the beta band in PD therapy [23, 2], careful frequency planning is a commonly used strategy, as previously discussed. However, this approach becomes less effective in the presence of large artifacts and constrains the flexibility of stimulation parameters, thereby limiting the potential for optimizing therapeutic efficacy. Digital filtering and template subtraction (TS) are widely used back-end methods. TS assumes artifact stability over time, averaging repeated instances to subtract from the signal [13, 35]. However, time-varying brain responses may alter artifact characteristics [24], limiting TS performance. To address this, Shrinkage and Manifold-based Artifact Removal using Template Adaptation (SMARTA) was introduced [24]. By leveraging manifold denoising and random matrix theory, SMARTA adapts templates for individual artifacts using the underlying geometry of stimulation-induced distortions and the high-dimensional nature of LFP noise. This enables precise artifact removal while preserving neural signals. Despite its strengths, SMARTA has two key limitations. First, it is computationally intensive: processing a 7-ms segment takes over 200 ms, making it unsuitable for real-time aDBS. Second, like TS, it does not resolve DC transient artifacts, which limits its use for short or repeated stimulation periods and impairs beta burst detection. To support real-time aDBS with an accurate beta burst detection, a more efficient and DC-transient-aware variant of SMARTA is needed for accurate and timely biomarker extraction.
@@
In this study, we refined SMARTA to address its limitations, resulting in SMARTA+, which preserves SMARTA’s core strengths while greatly improving computational efficiency and mitigating DC transient artifacts. A key enhancement involved replacing SMARTA’s most computationally intensive step, the k-nearest neighbors (KNN) search, with an approximate nearest neighbors (ANN) algorithm [4, 17, 18]. Unlike KNN, which performs exhaustive searches, ANN uses decision trees to identify similar artifacts with significantly reduced computational burden [4]. Additionally, we introduced a dimensionality reduction step based on wavelet transforms, which exploits the sparsity structure of stimulus artifacts [9]. These modifications enable the construction of a large stimulus artifact library, potentially from other subjects, allowing SMARTA+ to match each artifact with the most similar examples across datasets. This cross-subject learning improves artifact estimation and suppression, especially during stimulation onset, where DC transients alter artifact morphology (see Figure 1). Without a sufficiently diverse library, onset artifacts are particularly difficult to model. To further suppress DC transients, we incorporated a line-fitting algorithm that estimates and removes the slow-varying baseline shift at the start of each stimulation period. Together, these advances enable SMARTA+ to achieve effective artifact removal with a fraction of the original computational cost, advancing it toward real-time aDBS deployment.
@@
We evaluated the performance of SMARTA+ using aDBS signals recorded from PD patients and semi-real aDBS signals synthesized from these recordings. In addition to standard metrics, including normalized mean square error (NMSE) for quantifying LFP recovery across spectral bands, artifact residual (AR) for time-domain suppression, and spectral concentration (SC) for spectral contamination [24], we introduced a temporal event localization analysis to assess how accurately SMARTA+ recovers beta burst events. Specifically, we evaluated recall, precision, and F1-score to quantify the detection of beta burst onset and offset. This temporal event localization analysis, while previously applied in sEMG-based closed-loop DBS systems [30] for evaluating closed-loop DBS with post-stimulation control signals and sleep apnea detection from the photoplethysmogram signal [5], have not yet been used to evaluate aDBS algorithms that rely on real-time beta activity from LFPs. To our knowledge, SMARTA+ is the first algorithm explicitly designed to preserve beta burst timing. For benchmarking, we compared its performance against template subtraction [7], SMARTA [24] and blanking methods [1].
@@ 2. Materials
Patients undergoing bilateral implantation of STN-DBS for the treatment of advanced PD were recruited. The surgical procedures were described in [6]. LFPs were recorded one day after electrode implantation through the externalized electrode using Neuro Omega (Alpha Omega Engineering, Israel). All surgeries and recordings were conducted at Chang Gung Memorial Hospital, Linkou, Taoyuan, Taiwan, between 2023 and 2024 by the coauthors Dr. P. H. Tu and Dr. M. C. Yeap.
@@
The aDBS protocol followed the method in [23]. A second-order Butterworth band-pass filter (3-37 Hz) was first applied to the baseline LFP signal, the peak beta frequency (13-35 Hz) was identified, and a peak IIR filter with a Q factor of 3 centered at this frequency was applied three times to enhance beta activity. The resulting signal was then rectified and smoothed using a 400-ms moving average to extract beta amplitude. Stimulation was driven by a threshold set at the 75th percentile of the amplitude distribution. During the aDBS procedure, the online beta amplitude was calculated using the same method. Stimulation was initiated when the amplitude exceeded the predefined threshold and was turned off once it fell below that threshold. A minimum stimulation duration of 400 ms was enforced.
@@ 3. Proposed algorithm – SMARTA+
The SMARTA+ algorithm builds on the previously proposed SMARTA framework (Shrinkage and Manifold-based Artifact Removal using Template Adaptation) [24]. SMARTA+ introduces two key enhancements. First, to reduce computational burden, the original k-nearest neighbors (KNN) search is replaced with an ANN algorithm, and dimensionality is reduced via wavelet transform. These changes enable the construction of a larger and more diverse stimulus artifact library from available recordings, improving the accuracy of artifact removal. Second, a curve-fitting procedure is incorporated to better handle transient DC artifacts.
@@
Fig. 2 is the flowchart of SMARTA+, and the overall workflow is as follows. To address transient DC artifacts, a curve-fitting procedure is applied (Section 3.1) for all recordings. Prior to artifact removal, a stimulus artifact library is constructed from all available recordings (Section 3.2), potentially including data from other subjects. Then, recover each stimulus artifact applying the following algorithm, which is a modified SMARTA. For each recording we want to eliminate the stimulus artifacts, construct a decision forest from all recordings, where the wavelet transform is applied. For each target stimulus artifact, similar artifacts are retrieved from the library using the constructed decision forest and ANN (Section 3.3), where the leaf nodes at the end of the decision paths form a candidate template pool. Within this pool, the $K$ nearest neighbors are identified, with the optimal $K$ selected via greedy search (Section 3.4). The final artifact template is computed as the median of these neighbors and subtracted from the signal segment. The same curve-fitting step is applied again to further remove residual transient DC artifacts (Section 3.1), yielding the final artifact-free LFP. Below, we detail the algorithm step by step.
Assume the stimulation frequency is $f_{\mathrm{sti}}$, which is $130$ Hz in this study. Assume there are $M$ aDBS recordings that we can leverage and we want to recover the LFP of the $(M+1)$th aDBS recording.
@@ 3.1. Transient DC artifact removal
For each recording, the transient DC artifacts were removed prior to the segmentation of stimulus artifacts. Over each stimulation period, the trend that contains transient DC artifact is estimated by the following two-step procedure. The first step is applying a moving-average filter with a window length of 10 ms to eliminate large outliers in the signal and the second phase is fitting a sixth-order polynomial function to the smoothed signal. For non-stimulation periods, the trend was estimated and subtracted using the same procedures, except that a first-order polynomial was used for fitting.
@@ 3.2. Stimulus artifact library construction
To construct the stimulus artifact library, all available aDBS recordings were considered, and details are listed in the following two subsections.
@@ 3.2.1. Stimulus artifacts detection and segmentation
For the $m$-th recording, where $m=1,\ldots,M+1$, the peak-finding algorithm described in [24] was applied to obtain accurate timing of stimulus artifacts. Before identifying peaks, a third-order high-pass Butterworth filter with a cutoff frequency of 300 Hz was used to remove low-frequency noise. Additionally, a moving-average filter with a window length of 1 s was applied to estimate the trend, which was subsequently subtracted from the signal. The absolute value of the signal was than taken, followed by smoothing using a moving-average filter with a window length of 0.5 ms. A sample point was identified as a peak if it was greater than its neighboring points and exceeded a predefined threshold. This threshold was set as the 95th percentile of all sample points. Furthermore, the interval between detected peaks had to be greater than $1/f_{\mathrm{sti}}-\tau$, where $f_{\mathrm{sti}}$ is the stimulation frequency and $\tau$ is a bias-handling factor. In this study, $\tau$ was set to 0.5 ms. The detected peaks marked the locations of stimulus artifacts for segmentation. For aDBS signals, stimulation periods were further identified by detecting intervals between artifacts that were longer than the minimum period set in the aDBS algorithm.

### v2

@@ Abstract.
Background: Adaptive deep brain stimulation (aDBS) leverages symptom-related biomarkers to deliver personalized neuromodulation therapy, with the potential to improve treatment efficacy and reduce power consumption compared to conventional DBS. However, stimulation-induced signal contamination remains a major technical barrier to advance its clinical application.
New method: Existing artifact-removal strategies, both front-end and back-end, face trade-offs between artifact suppression and algorithmic flexibility. Among back-end algorithms, Shrinkage and Manifold-based Artifact Removal using Template Adaptation (SMARTA) has shown promising performance in mitigating stimulus artifacts with minimal distortion to local field potentials (LFPs), but its high computational demand and inability to handle transient direct current (DC) artifacts limits its use in real-time applications. To address this, we developed SMARTA+, a computationally efficient extension of SMARTA capable of suppressing both stimulus artifacts and DC transient artifacts while supporting flexible algorithmic design.
Results: We evaluated SMARTA+ using semi-real aDBS data and real data from Parkinson’s disease patients. It preserved the spectral and temporal structure of the underlying LFPs and demonstrated robustness across a variety of simulated stimulation protocols. Furthermore, the temporal event localization analysis showed that SMARTA+ can accurately determine the beta-burst events.
Comparison with existing methods: SMARTA+ outperformed template subtraction, pulse blanking, and transient blanking, and achieved performance comparable to or better than SMARTA while substantially reducing computation time.
Conclusions: By enhancing artifact suppression and improving computational efficiency, we show that SMARTA+ has the potential to advance real-time, closed-loop aDBS systems for neuromodulation therapies across diverse neurological disorders.
@@ 1. Introduction
Deep brain stimulation (DBS) is a widely used and effective therapy for advanced Parkinson’s disease (PD) [limousin1995DBS, benabid2003DBS, krack2003DBS, deuschl2006DBS]. By delivering electrical stimulation through implanted electrodes in the subthalamic nucleus (STN), DBS can alleviate motor impairments and improve patients’ quality of life. Additionally, local field potentials (LFPs) can be recorded from the same electrodes during stimulation, providing real-time insights into PD-related neural activity. Biomarkers derived from LFP enable adaptive deep brain stimulation (aDBS), where stimulation parameters are dynamically adjusted to respond to fluctuations in disease state [little2013aDBS, arlotti2018aDBS8hour]. The clinical benefits of aDBS have been demonstrated in PD, with improvements observed in motor symptoms such as tremor [velisar2019aDBS] and bradykinesia [little2016aDBS]. Beyond its application in PD, aDBS has also shown promise in treating essential tremor [yamamoto2013ET, opri2020ET] and epilepsy [kossoff2004aDBSinEpilepsy, stypulkowski2014aDBSinEpilepsy]. These studies suggest that aDBS may be broadly applicable to various movement disorders, offering a more efficient and patient-specific therapeutic strategy.
@@
When the frequency band of interest is known in advance, careful selection of stimulation frequency and sampling rate (known as frequency planning or interleaving) [thenaisie2021percept] is often considered to mitigate aliasing, especially in preserving the beta band for Parkinson’s disease (PD) treatment [little2013aDBS, arlotti2018aDBS8hour]. However, this approach has limitations, since the dynamic morphology of stimulation artifacts over time can still degrade signal quality via spectral leakage and aliasing, particularly in the presence of large artifacts, and thereby limit the potential for optimizing therapeutic efficacy. Moreover, even if beta band activity (13-35 Hz) can be well preserved, extending protection to other relevant frequency bands remains challenging, which constrains the flexibility of stimulation parameters and its generalization to other applications. Gamma (60-90 Hz) [swann2018aDBSgamma, lofredi2018aDBSgamma] and alpha band activity (4-8 Hz) [marceglia2011alpha] have both been proposed as potential biomarkers for aDBS in PD, and high-frequency oscillations (HFOs) above 200 Hz have been shown to be associated with tremors in PD [hirschmann2016HFO]. Very high-frequency oscillation (VHFO) in the 1000-2500 Hz range [usui2015VHFO] have shown utility in epilepsy and may also hold relevance for PD. Furthermore, accurately removing stimulus artifacts alone may not suffice. The abrupt switching of stimulation often induces direct current (DC) transient artifacts at stimulation onset and offset, which can further contaminate biomarkers [anso2022artifactRemoval]. Together, these artifact-related challenges not only compromise the fidelity of LFP-based biomarkers but also impair aDBS control algorithms by introducing erroneous feedback and triggering unintended stimulation [neumann2023aDBSreview]. These observations underscore the need for robust and flexible artifact removal algorithms capable of preserving a broad spectrum of physiologically meaningful signals.
@@
In general, a basic front-end approach is common-mode rejection, where two sensing contacts adjacent to the stimulation electrode are used to obtain bipolar LFPs by signal subtraction [little2013aDBS, arlotti2018aDBS8hour]. Ideally, when contacts are symmetrically positioned around the stimulation site, artifacts cancel out. However, impedance mismatches and tissue asymmetries often leave residual artifacts and DC transient distortions [anso2022artifactRemoval]. To further reduce artifacts, blanking techniques can be applied at the front-end [stanslaski2018aDBSdevice] or back-end [heffer2008interpolation], temporarily disabling signal acquisition during stimulation pulses. While effective at suppressing high-amplitude artifacts, blanking also removes underlying neural signals and fails to address transient DC artifacts that persist beyond the stimulation period. Digital filtering and template subtraction (TS) are widely used back-end methods. TS assumes artifact stability over time, averaging repeated instances to subtract from the signal [hashimoto2002templateSub, sun2016MVS]. However, time-varying brain responses may alter artifact characteristics [liu2024], limiting TS performance. To address this, Shrinkage and Manifold-based Artifact Removal using Template Adaptation (SMARTA) was introduced [liu2024]. By leveraging manifold denoising and random matrix theory, SMARTA adapts templates for individual artifacts using the underlying geometry of stimulation-induced distortions and the high-dimensional nature of LFP noise. This enables precise artifact removal while preserving neural signals, outperforming existing algorithms.
Despite its strengths, SMARTA has two key limitations. First, it is computationally intensive: processing a 7-ms segment under 22 kHz sampling rate takes over 200 ms, making it unsuitable for real-time aDBS. Second, like TS, it does not resolve DC transient artifacts, which limits its use for short or repeated stimulation periods and impairs beta burst detection. To support real-time aDBS with an accurate beta burst detection, a more efficient and DC-transient-aware variant of SMARTA is needed for accurate and timely biomarker extraction. Although the theoretical framework of SMARTA allows for signals with heterogeneous sampling rates and stimulation frequencies, a systematic evaluation of this capability remains lacking. Moreover, to our knowledge, a reliable detection of dynamic events, for instance, the onset and offset of beta bursts in PD patients, is less discussed in the literature.
@@
In this study, we refined SMARTA to address its limitations, resulting in SMARTA+, which preserves SMARTA’s core strengths while greatly improving computational efficiency, mitigating DC transient artifacts and reliably detecting beta bursts. A key enhancement involved replacing SMARTA’s most computationally intensive step, the k-nearest neighbors (KNN) search, with an approximate nearest neighbors (ANN) algorithm [ANNOY, jones2011RANN, jones2013RANN]. Unlike KNN, which performs exhaustive searches, ANN uses decision trees to identify similar artifacts with significantly reduced computational burden [ANNOY]. Additionally, we introduced a dimensionality reduction step based on wavelet transforms, which exploits the sparsity structure of stimulus artifacts [donoho1995wavelet]. These modifications enable the construction of a large stimulus artifact library, potentially from other subjects, allowing SMARTA+ to match each artifact with the most similar examples across datasets. This cross-subject learning improves artifact estimation and suppression, especially during stimulation onset, where DC transients alter artifact morphology (see Figure 1). Without a sufficiently diverse library, onset artifacts are particularly difficult to model. To further suppress DC transients, we incorporated a line-fitting algorithm that estimates and removes the slow-varying baseline shift at the start of each stimulation period. Together, these advances enable SMARTA+ to achieve effective artifact removal with a fraction of the original computational cost, advancing it toward real-time aDBS deployment.
@@
We evaluated the performance of SMARTA+ using aDBS signals recorded from PD patients and semi-real aDBS signals synthesized from these recordings. For benchmarking, we compared its performance against template subtraction [culaclii2016ts], SMARTA [liu2024] and blanking methods [anso2022artifactRemoval]. In addition to standard evaluation metrics, including the normalized mean square error (NMSE), artifact residual (AR), and spectral concentration (SC), which respectively quantify LFP recovery across spectral bands, in the time domain, and in the spectral domain [liu2024], we introduced a temporal event localization analysis to assess how accurately SMARTA+ recovers beta burst events. Specifically, we evaluated recall, precision, and F1-score to quantify the detection of beta burst onset and offset. This temporal event localization analysis, while previously applied in sEMG-based closed-loop DBS systems [shukla2012] for evaluating closed-loop DBS with post-stimulation control signals and sleep apnea detection from the photoplethysmogram signal [chen2025validation], have not yet been used to evaluate aDBS algorithms that rely on real-time beta activity from LFPs. To our knowledge, SMARTA+ is the first algorithm explicitly designed to preserve beta burst timing.
@@ 2. Materials
Patients with advanced Parkinson’s disease who underwent bilateral STN-DBS implantation were included through secondary use of clinical data obtained from prior IRB-approved studies. The surgical procedures have been described previously [chen2021YCC]. Local field potentials (LFPs) were recorded one day after electrode implantation using externalized leads and the Neuro Omega system (Alpha Omega Engineering, Israel). All surgeries and recordings were performed at Chang Gung Memorial Hospital, Linkou, Taoyuan, Taiwan, between 2023 and 2024 by the coauthors Dr. P. H. Tu and Dr. M. C. Yeap.
@@
The aDBS protocol followed the method of [little2013aDBS]. A second-order Butterworth band-pass filter (3-37 Hz) was first applied to the baseline LFP signal. The peak beta frequency (13-35 Hz) was identified, and a peak IIR filter with a Q factor of 3 centered at this frequency was applied three times to enhance beta activity. The resulting signal was then rectified and smoothed using a 400-ms moving average to extract beta amplitude. Stimulation was driven by a threshold set at the 75th percentile of the amplitude distribution. During the aDBS procedure, the online beta amplitude was calculated using the same method. Stimulation was initiated when the amplitude exceeded the predefined threshold and was turned off once it fell below that threshold. A minimum stimulation duration of 400 ms was enforced.
@@ 3. Proposed algorithm – SMARTA+
The SMARTA+ algorithm builds on the previously proposed SMARTA (Shrinkage and Manifold-based Artifact Removal using Template Adaptation) framework [liu2024]. SMARTA+ introduces two key enhancements. First, to reduce computational burden, the original KNN search is replaced with an ANN algorithm, and dimensionality is reduced via wavelet transform. These updates enable subject-adaptive, online construction of a larger and more diverse stimulus artifact library, improving artifact removal accuracy and enabling real-time implementation. Second, a curve-fitting procedure is incorporated to better handle transient DC artifacts.
@@
The flowchart of SMARTA+ is shown in Fig. 2. The overall workflow is as follows. A curve-fitting procedure is applied for all recordings to address transient DC artifacts (Section 3.1). Prior to removing artifacts from a new recording, indexed by $n+1$, a stimulus artifact library is constructed from all available $n$ recordings (Section 3.2), potentially including data from different subjects. With the library, each stimulus artifact in the $(n+1)$th recording is recovered by applying the following algorithm, which is a modification of SMARTA. First, construct a decision forest from the library, where the wavelet transform is applied. For each target stimulus artifact in the $(n+1)$th recording, similar artifacts are retrieved from the library using the constructed decision forest and ANN (Section 3.3), where the leaf nodes at the end of the decision paths form a candidate template pool. Within this pool, the $K$ nearest neighbors are identified, with the optimal $K$ selected via greedy search (Section 3.4). The final artifact template is computed as the median of these neighbors and subtracted from the signal segment. After removing stimulation artifacts, the same curve-fitting step is applied again to improve the residual transient DC artifact removal (Section 3.1), yielding the final artifact-free LFP.
Below, we detail the algorithm step by step. Denote the stimulation frequency by $f_{\mathrm{sti}}$, which is typically 130 Hz in DBS therapy for PD. Assume there are $M$ aDBS recordings available and we aim to recover the LFP of the $(M+1)$th aDBS recording.
@@ 3.1. Transient DC artifact removal
For each recording, transient DC artifacts were removed prior to the segmentation of stimulus artifacts. Over each stimulation period, the trend containing the transient DC artifact is estimated by the following two-step procedure. The first step is to apply a moving-average filter with a window length of 10 ms to eliminate large outliers in the signal, and the second step is to fit a sixth-order polynomial function to the smoothed signal. For non-stimulation periods, the trend was estimated and subtracted using the same procedures, except that a first-order polynomial was used for fitting.
@@ 3.2. Stimulus artifact library construction
The library is constructed in two stages: artifact detection and segmentation, followed by ANN preparation, including data cleanup and dimension reduction. In real-time use, the target recording under analysis is excluded.
@@ 3.2.1. Stimulus artifacts detection and segmentation
For the $m$-th recording, where $m=1,\ldots,M+1$, the peak-finding algorithm described in [liu2024] was applied to obtain accurate timing of stimulus artifacts. Before identifying peaks, a third-order high-pass Butterworth filter with a cutoff frequency of 300 Hz was used to remove low-frequency noise. Additionally, a moving-average filter with a window length of 1 s was applied to estimate the trend, which was subsequently subtracted from the signal. The absolute value of the signal was then taken, followed by smoothing using a moving-average filter with a window length of 0.5 ms. A sample point was identified as a peak if it was greater than its neighboring points and exceeded a predefined threshold. This threshold was set as the 95th percentile of all sample points. Furthermore, the interval between detected peaks had to be greater than $1/f_{\mathrm{sti}}-\tau$, where $f_{\mathrm{sti}}$ is the stimulation frequency and $\tau$ is a bias-handling factor. In this study, $\tau$ was set to 0.5 ms. The detected peaks marked the locations of stimulus artifacts for segmentation. For aDBS signals, stimulation periods were further identified by detecting intervals between artifacts that were longer than the minimum period set in the aDBS algorithm.